### v1

# Can LLMs Track Their Output Length? A Dynamic Feedback Mechanism for Precise Length Regulation

Abstract
Precisely controlling the length of generated text is a common requirement in real-world applications. However, despite significant advancements in following human instructions, Large Language Models (LLMs) still struggle with this task. In this work, we demonstrate that LLMs often fail to accurately measure input text length, leading to poor adherence to length constraints. To address this issue, we propose a novel length regulation approach that incorporates dynamic length feedback during generation, enabling adaptive adjustments to meet target lengths. Experiments on summarization and biography tasks show our training-free approach significantly improves precision in achieving target token, word, or sentence counts without compromising quality. Additionally, we demonstrate that further supervised fine-tuning allows our method to generalize effectively to broader text-generation tasks.
Can LLMs Track Their Output Length?
A Dynamic Feedback Mechanism for Precise Length Regulation
Meiman Xiao1††thanks: These authors contributed equally., Ante Wang111footnotemark: 1, Qingguo Hu111footnotemark: 1, Zhongjian Miao3, Huangjun Shen1 Longyue Wang2, Weihua Luo2, Jinsong Su1††thanks: Corresponding author. 1School of Informatics, Xiamen University, China 2Alibaba International Digital Commerce Group 3Li Auto Inc. {xiaomeiman,wangante,huqingguo}@stu.xmu.edu.cn jssu@xmu.edu.cn
1 Introduction
The use of products built upon Large Language Models (LLMs, achiam2023gpt; team2023gemini; yang2025qwen3) to process various text generation tasks has grown prevalent in daily human study and work. Benefiting from pretraining on massive corpora and subsequent post-training, such as Supervised Fine-Tuning (SFT, taori2023alpaca; chiang2023vicuna) and Reinforcement Learning (RL, ouyang2022training; bai2022training), current LLMs demonstrate strong text comprehension and instruction-following capabilities, enabling them to effectively handle diverse user requests.
Despite their significant progress, it remains challenging for LLMs to precisely control their outputs according to specific requirements. For instance, when given a query such as “Please write a 500-word composition,” current LLMs often struggle with strictly adhering to the length constraint. To address this issue, previous research jie_prompt-based_2024; yuan2024following; li2024ruler; butcher_precise_2025 has trained LLMs to produce outputs that adhere to specified length requirements through SFT or advanced RL algorithms. However, these approaches often demand massive training data to cover a wide range of length preferences, and may degrade model performance after intensive post-training. More critically, these techniques are typically tailored to a single type of length constraint, hindering generalization to different types of length control (e.g., tokens, words, or sentences).
Intuitively, an LLM should track its output length to precisely control the length of its generated content. To explore this issue further, we first conduct a preliminary study to investigate: Do LLMs know how long their generated texts are? Our results indicate that while LLMs can accurately measure the length of short texts, this ability declines significantly as text length increases. This deficiency mirrors human tendencies: when writing, people typically focus on content creation and planning rather than precisely tracking exact length. However, unlike LLMs, humans can leverage tools such as text editors to obtain external feedback for monitoring counts, which frees them from manual estimation and allows them to use this feedback to adjust their strategies for subsequent content.
Inspired by these findings, we explore the use of length feedback to help LLMs track and dynamically adjust their output length. Previous research markergen successfully incorporated word-level length guidance through a three-stage pipeline, inserting length feedback after tokens with an exponential decay strategy. However, this method incurs high inference overhead, and its iterative process can lead to error propagation, potentially harming response quality. In contrast, our approach has LLMs generate text in one pass but intermittently interrupts the generation process to insert feedback. To ensure the LLM understands the feedback, we enrich the prompt with explanations of its meaning and usage. To minimize disruption, we insert feedback only at sentence boundaries, thereby preserving semantic coherence. More importantly, our method offers flexible adaptation to different length constraints by simply adjusting the feedback style and input prompt. A comparison with the conventional prompt-based approach is shown in Figure 1.
We first validate our method on summarization huang-etal-2021-efficient and biography generation min2023factscore tasks. Without specialized fine-tuning, our method demonstrates superior length control across different constraint types when applied to LLaMA-3.1-8B-Instruct dubey2024llama and Qwen3 yang2025qwen3 family models. Evaluations based on DeepSeek-V3 deepseekai2024deepseekv3technicalreport indicate that the text quality remains robust while meeting these constraints.
To assess our method’s effectiveness on general-domain tasks, we further conduct experiments on ELI5 eli5, a benchmark for long-form question answering. While our training-free approach can better meet length requirements, it may compromise quality for tasks requiring answers within a strict length range. However, by applying SFT only to LLM-generated responses and without any data filtering, our method effectively maintains performance while enhancing length control. Further analysis shows that our method achieves at least 4 greater training efficiency than baselines without feedback incorporation, avoiding performance degradation from excessive fine-tuning.
2 Preliminary Study
Generating text with a desired length inherently requires the model to accurately track its output length, yet the LLM’s ability to estimate length remains unclear. In this section, we design experiments to investigate this question: Can LLMs accurately estimate the length of their generated texts? Specifically, we first collect responses of varying lengths from several mainstream LLMs and evaluate their length-tracking ability by prompting them to estimate the lengths of their own outputs (§2.1). We then present the evaluation results and highlight the key findings (§2.2).
2.1 Experimental Setup
Data Preparation
Following prior work jie_prompt-based_2024; butcher_precise_2025, we conduct experiments on the summarization task, where length constraints are commonly specified. We use the GovReport dataset huang-etal-2021-efficient and evaluate the following LLMs: LLaMA-3.1-8B-Instruct dubey2024llama, as well as Qwen3-4B and Qwen3-8B yang2025qwen3, chosen for their strong reasoning capabilities on challenging tasks. To obtain responses of varying lengths, we prompt the LLMs to generate summaries for each instance with a randomly selected target length between 100 and 400 tokens. In total, 1,000 generated summaries are collected.
Length Estimation
To evaluate the LLMs’ ability to track output length, we feed each generated summary back into the same model and prompt it to estimate its length. Specifically, we use the prompt in Figure 16 to ask the LLM for measuring the output length in tokens. We use the Mean Absolute Error (MAE) between the LLM-estimated and actual generated length to quantify the LLMs’ prediction errors in counting output length. Additionally, we report the MAE between the LLM-generated and user-specified lengths to provide a more direct assessment of the LLMs’ length controlling capability.
2.2 Results and Analysis
Results in Figure 2 offer several key observations.
LLMs perform worse when counting longer texts
Considering the MAE between the LLM-estimated and generated text length, we find a strong positive correlation between prediction errors and text length across all three LLMs. This suggests that as text length increases, LLMs’ estimations deviate more significantly from the true length, revealing a fundamental limitation in their length-tracking capability. These trends also persist across other types of length constraints (e.g., words and sentences). Detailed results are provided in the Appendix A.
Length control is closely correlated with length estimation
We observe the same trend when comparing the prediction errors of LLM-generated and LLM-estimated lengths against the corresponding target length. For the Qwen3 series models, those that achieve more accurate length estimation also tend to perform better in length control. This conclusion holds across different types of length constraints. For example, when estimating output length, LLaMA-3.1-8B-Instruct performs significantly better at word counting than token counting, leading to greater compliance with word-length constraints over token-based ones in generation. These results underscore the importance of length tracking in length control.
3 Methodology
In our preliminary study, we observe that existing LLMs struggle to accurately track their output length, resulting in poor control over text generation. To overcome this challenge, we consider how humans handle similar length control problems during writing. Rather than relying on manual estimation, humans typically use external tools such as text editors to precisely monitor the length of their ongoing work. Abandoning inefficient multi-stages used in Markergen, we propose a novel length regulation approach that integrates length feedback from an external counter at the end of sentences to help LLMs track the length of their generated output more effectively and efficiently.
As shown in Figure 1(b), we intermittently interrupt the LLM’s generation to insert length feedback. Formally, given an unfinished output , we append a feedback string (e.g., “<used_tokens=>”) before resuming generation, where is a predefined length calculation function. Depending on the desired length constraint type (e.g., tokens, words, sentences), can be easily implemented using the model’s tokenizer or external tools like NLTK loper2002nltk. Specifically, during generation, we detect sentence boundaries and invoke the length-calculation tool based on the specified constraint type. This design minimizes disruption to the natural flow of generation while maintaining semantic coherence.
To ensure the LLM correctly interprets this length feedback, we augment the prompt with instructions explaining its meaning and usage, as illustrated below:
Leveraging the strong comprehension capabilities of LLMs, we find that they can correctly interpret the feedback string without disrupting the natural generation fluency. In experiments, we further construct a feedback-augmented dataset and demonstrate that SFT enables LLMs to actively use the tool and further enhances response quality.
4 Experiments
In this section, we first detail the experimental setup (§4.1). Next, we validate our training-free approach on specific tasks (§4.2) and finally extend our method to the general domain via SFT (§4.3).
4.1 Setup
Datasets
We evaluate our approach across three different models on two tasks suited for length constraints, as well as a general-domain long-form generation task, to assess method robustness.
GovReport huang-etal-2021-efficient: A long-document summarization dataset of 973 U.S. government reports with expert-written abstractive summaries for testing. Unlike conventional short-form summarization datasets cohan-etal-2018-discourse; kornilova-eidelman-2019-billsum; sharma-etal-2019-bigpatent, GovReport allows better evaluation across various target lengths.
Biographies min2023factscore: A biography generation dataset of 183 individuals for testing, originally proposed to evaluate the factuality of LLM-generated text. Unlike summarization, this task relies on LLMs to generate content based on their parametric knowledge.
ELI5 eli5: A general-domain long-form question-answering dataset. We use a publicly available dataset111https://huggingface.co/datasets/sentence-transformers/eli5, randomly sampling 10,240 instances for training and 500 for testing.
| Token | Word | Sentence | |||||||
| MAE | PM | Qual. | MAE | PM | Qual. | MAE | PM | Qual. | |
| GovReport | |||||||||
| Qwen3-4B | 58.19 | 11.94 | 4.94 | 70.74 | 8.31 | 4.95 | 15.85 | 16.36 | 4.74 |
| ICL | 59.07 | 13.30 | 4.77 | 104.20 | 3.88 | 4.80 | 8.20 | 18.41 | 4.53 |
| Feedback | 61.30 | 21.35 | 4.64 | 31.96 | 38.53 | 4.84 | 1.38 | 47.83 | 4.62 |
| ICLFeedback | 26.33 | 34.46 | 4.71 | 26.63 | 39.91 | 4.83 | 0.51 | 58.40 | 4.67 |
| Qwen3-8B | 52.04 | 15.20 | 4.96 | 60.61 | 6.06 | 4.95 | 3.83 | 35.59 | 4.91 |
| ICL | 59.51 | 11.16 | 4.76 | 141.20 | 0.16 | 4.92 | 4.09 | 27.98 | 4.89 |
| Feedback | 22.81 | 58.24 | 4.84 | 21.23 | 47.50 | 4.93 | 0.57 | 79.47 | 4.82 |
| ICLFeedback | 17.40 | 50.55 | 4.76 | 51.17 | 17.29 | 4.89 | 0.10 | 94.81 | 4.88 |
| Markergen | 44.26 | 37.27 | 4.29 | 42.19 | 20.91 | 4.62 | 1.87 | 38.61 | 4.74 |
| LLaMA-3.1-8B-Instruct | 108.65 | 5.93 | 4.78 | 32.83 | 14.32 | 4.81 | 2.94 | 46.64 | 4.58 |
| ICL | 118.27 | 3.31 | 4.70 | 50.60 | 8.66 | 4.79 | 3.32 | 37.77 | 4.54 |
| Feedback | 33.88 | 41.30 | 4.64 | 16.22 | 43.17 | 4.72 | 0.25 | 80.10 | 4.63 |
| ICLFeedback | 25.94 | 38.80 | 4.53 | 22.93 | 35.69 | 4.72 | 0.22 | 82.06 | 4.54 |
| Markergen | 41.45 | 56.67 | 3.11 | 34.20 | 54.50 | 3.54 | 3.36 | 31.22 | 3.66 |
| Biographies | |||||||||
| Qwen3-4B | 80.46 | 8.24 | 2.94 | 27.66 | 29.82 | 3.04 | 10.94 | 20.58 | 2.80 |
| ICL | 25.07 | 29.01 | 2.94 | 29.41 | 30.45 | 2.98 | 3.36 | 21.68 | 2.62 |
| Feedback | 93.36 | 35.68 | 2.82 | 19.37 | 51.99 | 2.99 | 0.28 | 81.66 | 2.93 |
| ICLFeedback | 14.49 | 59.41 | 2.98 | 7.19 | 79.31 | 3.02 | 0.17 | 82.60 | 2.78 |
| Qwen3-8B | 78.85 | 8.74 | 3.14 | 55.11 | 8.59 | 3.12 | 2.67 | 26.44 | 3.11 |
| ICL | 27.70 | 28.98 | 3.11 | 38.86 | 16.25 | 3.09 | 1.25 | 50.82 | 3.07 |
| Feedback | 22.78 | 66.01 | 3.12 | 9.22 | 73.66 | 3.14 | 0.05 | 97.91 | 3.09 |
| ICLFeedback | 11.22 | 70.49 | 3.13 | 12.75 | 53.16 | 3.12 | 0.01 | 98.54 | 3.06 |
| Markergen | 42.91 | 44.40 | 2.49 | 26.79 | 40.20 | 2.94 | 2.08 | 33.26 | 3.02 |
| LLaMA-3.1-8B-Instruct | 97.64 | 6.98 | 3.12 | 51.72 | 10.23 | 3.22 | 4.86 | 24.46 | 3.02 |
| ICL | 41.81 | 21.16 | 3.24 | 28.69 | 20.04 | 3.33 | 6.12 | 17.12 | 3.12 |
| Feedback | 36.90 | 39.04 | 3.34 | 21.96 | 45.15 | 3.32 | 0.77 | 80.20 | 3.24 |
| ICLFeedback | 24.11 | 47.54 | 3.38 | 11.61 | 61.92 | 3.36 | 0.19 | 92.53 | 3.22 |
| Markergen | 69.26 | 53.83 | 2.91 | 37.16 | 53.94 | 2.97 | 2.68 | 29.55 | 2.90 |
Models and Hyperparameters
We use the same models as in the preliminary study (§2): Qwen3-4B, Qwen3-8B, and LLaMA-3.1-8B-Instruct. For all experiments, we conduct inference using the vLLM engine kwon2023efficient, which provides high-throughput serving for LLMs. We employ nucleus sampling with temperature set to 0.8 and top-p value of 0.95 for text generation. We define target lengths across three linguistic granularities: token-level (100 to 400 tokens, in increments of 50), word-level (100 to 400 words, in increments of 50), and sentence-level (5 to 30 sentences, in increments of 5). To quantify generation length across different linguistic units, we use each model’s native tokenizer for token-level counts and the NLTK loper2002nltk for word- and sentence-level measurements.
Evaluation Metrics
Following prior studies butcher_precise_2025; li2024ruler, we evaluate length‐generation error and text quality at these target lengths using the metrics defined below.
Mean Absolute Error (MAE): This metric computes the average absolute difference between the predicted length and the ground-truth user-specified length across all samples:
(1) -
Precise Match (PM): This metric calculates the fraction of predictions whose absolute error falls within a predefined tolerance threshold :
(2) where denotes the indicator function. In our experiments, we set for token and word counts, and for sentence-level measurements.
Quality Score (Qual.): We adopt the LLM-as-a-Judge paradigm zheng2023judgingllmasajudgemtbenchchatbot, using DeepSeek-V3222DeepSeek-V3-0324 across all datasets for more accurate assessments. Specifically, for GovReport, we follow liu-etal-2023-g to evaluate summaries on three dimensions: coherence, consistency, and relevance. For Biographies, we assess coherence and factual accuracy. For ELI5, we follow zheng2023judgingllmasajudgemtbenchchatbot by focusing on correctness and helpfulness. To manage costs, for each linguistic unit and specified length, we randomly select 100 instances for summary or biography evaluation and 200 instances for general-domain text generation. To ensure reliability, we further conduct a human evaluation (Appendix B), which validates strong consistency between human and model judgments. The Quality Score is reported as the average of the relevant metrics for each dataset.
4.2 Results on Training-Free Length Control
Given the high cost of fine-tuning LLMs, we first compare training-free approaches. In addition to conventional prompting with target length constraints (Baseline), we investigate the effectiveness of in-context learning (ICL), which exhibits strong performance across various tasks without task-specific fine-tuning wies2023learnability; dong2024survey. To implement this, we collect instances of varying lengths and construct a one-shot demonstration for each instance. Besides, we implement Markgen markergen as a competitive baseline, given that its official code is not publicly available.
Results on GovReport and Biographies are in Table 1. Our key findings are as follows:
Incorporating length feedback effectively improves length control while maintaining text quality
Comparisons of the Baseline and Feedback demonstrate consistent and significant improvements in both MAE and PM across all datasets, linguistic units, and LLMs. To further illustrate this, Figure 3 presents violin plots comparing the two methods with Qwen3-8B on the GovReport dataset at various target lengths. The results show that Feedback yields significantly tighter length distributions than Baseline across all granularities, with a higher density around the target lengths. This confirms the effectiveness of our method in enhancing length control. Additionally, most Qual. scores remain competitive across both datasets for different LLMs, suggesting that text quality is effectively preserved in our method. For Markergen, we also observe a significant improvement in length control. However, our method achieves relatively better results without expensive multi-stage inference. In terms of efficiency, our method is approximately 5 more efficient than Markergen in token consumption. Furthermore, Markergen exhibits noticeably poorer quality scores than the Baseline. A closer examination suggests that the inserted feedback within the body of a sentence can disrupt the content fluency, and LLMs may not adequately follow the prompt at each stage, leading to an error propagation issue.
Generating with length feedback benefits more from ICL than vanilla prompting
We observe that ICL has an unstable impact on the Baseline. Generally, it degrades the performance on GovReport, while bringing some benefits to Biographies. In contrast, for ICL+Feedback, we observe a more consistent improvement. This is because our method enables better leveraging of the input demonstration to learn how to use feedback information for controlling the generation length. However, we still observe some obvious decline when using Qwen3-8B, which may be due to its capability of leveraging additional contexts. Nevertheless, these findings suggest that our method can be further improved even with simple ICL, without requiring specific fine-tuning.
Stronger models gain more from incorporating length feedback
Generally, larger models possess stronger language understanding capabilities, enabling them to more effectively leverage length feedback to adjust subsequent outputs. When comparing Qwen-4B and Qwen-8B, we observe that larger models consistently achieve greater improvements in MAE and PM scores while typically experiencing only minor negative effects on text quality when incorporating length feedback. This suggests that our method can achieve better performance when integrated into LLM products, which generally have stronger baseline capabilities.
4.3 Results on Training-Based Length Control
Though incorporating length feedback can effectively improve length control, we notice that it sometimes blindly reaches the target length, ignoring the response quality for general-domain instances from ELI5. To address this issue, we further investigate SFT for LLMs to balance the length and quality goals.
| Token | Word | Sentence | |||||||
| MAE | PM | Qual. | MAE | PM | Qual. | MAE | PM | Qual. | |
| Qwen3-8B | 59.84 | 9.35 | 7.83 | 46.09 | 14.71 | 8.06 | 1.82 | 48.67 | 7.57 |
| ICL | 30.64 | 19.65 | 7.80 | 33.27 | 20.86 | 7.89 | 2.74 | 34.57 | 7.71 |
| Feedback | 60.99 | 37.09 | 6.78 | 20.86 | 56.29 | 7.50 | 0.12 | 91.15 | 7.51 |
| ICLFeedback | 10.33 | 68.65 | 7.83 | 20.65 | 39.13 | 7.85 | 0.08 | 92.80 | 7.69 |
| SFT | 23.72 | 30.59 | 8.14 | 27.68 | 24.86 | 8.08 | 2.21 | 19.18 | 8.18 |
| SFTFeedback | 16.41 | 47.90 | 8.01 | 14.10 | 53.94 | 8.07 | 0.40 | 78.60 | 8.11 |
| RULER | – | – | – | 20.80 | 34.55 | 6.70 | – | – | – |
| LLaMA-3.1-8B-Instruct | 70.03 | 9.05 | 7.44 | 39.58 | 9.80 | 7.45 | 3.52 | 33.45 | 7.26 |
| ICL | 53.83 | 9.45 | 7.43 | 35.37 | 15.03 | 7.52 | 4.05 | 27.84 | 7.31 |
| Feedback | 24.48 | 37.94 | 7.38 | 15.55 | 44.47 | 7.33 | 0.11 | 95.75 | 7.05 |
| ICLFeedback | 25.28 | 39.18 | 7.25 | 14.15 | 50.31 | 7.26 | 0.30 | 87.13 | 7.18 |
| SFT | 31.91 | 18.60 | 7.49 | 29.24 | 19.06 | 7.46 | 3.43 | 11.12 | 7.50 |
| SFTFeedback | 18.39 | 46.41 | 7.28 | 15.40 | 52.53 | 7.40 | 1.54 | 44.40 | 7.45 |
| RULER | – | – | – | 27.31 | 29.46 | 6.10 | – | – | – |
For each instance in our prepared ELI5 training set, we assign a randomly generated target length to create a length requirement. We then provide both the user question and this requirement to the LLM for response generation. Due to the original LLM’s limited ability to follow length constraints, the sampled responses often deviate from the specified target length. To address this, we adjust the target length in the prompt based on the response length, as specified in Algorithm 1. For simplicity, we do not filter any response based on its quality due to the costly evaluation. Nevertheless, we show that this can effectively preserve the performance. In the prompt-based SFT, we use the original model responses without modification. For the feedback-based setting, we first segment responses into sentences, then append length feedback signals after each sentence. In addition, we implement RULER li2024ruler, a method that employs a meta-token for word-level length control. We utilize the official code333https://github.com/Geaming2002/Ruler for this implementation. To ensure a fair comparison, we train RULER using the same dataset employed for prompt-based SFT. Training details are reported in Appendix E.
Results on ELI5 are presented in Table 2. The following conclusions can be drawn:
Incorporating length feedback without fine-tuning sometimes degrades performance on the general domain
Mirroring the results on specific tasks, incorporating length feedback without fine-tuning continues to yield consistent improvements in open-domain question answering. However, we observe a decline in text quality for certain ELI5 settings, particularly with Qwen3-8B. This suggests the added feedback can sometimes mislead standard generation. We find, however, that ICL can partly alleviate this problem for Qwen3-8B, underscoring the need for further optimization to preserve output quality.
SFT effectively balances length control and generation quality
Applying SFT generally improves length control performance for both the Baseline and Feedback at the token and word levels, echoing the effectiveness of prior approaches. RULER achieves better length control performance than SFT, however, at the cost of response quality, possibly due to the introduced the meta-token, influncing the orignal generation. For SFT+Feedback, results demonstrate notable improvement in Qual. scores, performing on par with the Baseline. However, we observe that its sentence-level length control is poorer than that of Feedback. This may be because it is harder to balance the objectives of sentence number and generation quality on this dataset. Nevertheless, our method still substantially outperforms the Baseline.
Length feedback boosts post-training efficiency
Intuitively, length feedback frees LLMs from the burden of tracking output length, allowing them to achieve greater training efficacy in length control. To demonstrate this, we analyze the training dynamics of Qwen3-8B at token-level granularity in Figure 5. Results show that SFT+Feedback is at least 4 more efficient than SFT, achieving competitive performance in just 32 steps. Moreover, we observe that SFT+Feedback has not yet fully converged, suggesting that further scaling the training set could lead to even better performance.
Feedback-based SFT exhibits stronger generalization across length units and tasks
We apply trained models on character-level granularity. Results in Figure 6 demonstrate that SFT+Feedback consistently outperforms the baseline, with a higher density around the target length in each violin plot. Furthermore, we assess cross-task generalization on TruthfulQA lin-etal-2022-truthfulqa. As shown in Appendix F, our model achieves better cross-task generalization. We believe enriching the training set with more diverse length constraints and tasks could further improve performance. We leave this exploration for future work.
5 Related Work
5.1 Length Control for Text Generation
Controlling output length is crucial to meet diverse user requirements. Early studies addressed this challenge by designing early-stopping decoding strategies rush2015neural or incorporating specialized length-aware module liu2018controlling; takase2019positional; yu2021lenatten to guide generation. With the emergence of LLMs achiam2023gpt, their advanced instruction-following capability enables length control through direct prompting markergen. However, existing methods still rely on extensive post-training, such as SFT yuan2024following; li2024ruler; butcher_precise_2025 or RL singhal2023long; jie_prompt-based_2024, to achieve precise length control. In contrast, our method accomplishes this goal by incorporating length feedback, which is either training-free or training-efficient. This approach enables strong generalization under diverse length constraints.
5.2 Feedback-Guided Text Generation
Since LLMs are essentially probabilistic models, they often struggle with tasks involving complex mathematical reasoning zhou2022least or up-to-date factual knowledge luo2023systematic. A straightforward way to overcome these limitations is to equip LLMs with the ability to use external tools parisi2022talm; schick2023toolformer, such as search engines, calculators, or compilers. Using external feedback from these tools, LLMs’ capabilities can be substantially extended, allowing them to focus better on tasks they inherently excel at. Markergen markergen was the first to integrate length feedback directly into the generation process. However, its multi-stage framework introduces considerable complexity and inefficiency. In contrast, empirical results demonstrate that our method achieves better length control while maintaining text quality in a single pass. This streamlined process naturally allows further optimization through training.
6 Conclusion
In this work, we propose incorporating external length feedback at the end of sentences during LLM generation. To enable models to internalize this mechanism, we further construct a dataset enriched with length feedback and perform SFT, promoting generalization across diverse tasks and length specifications. We evaluate our approach on both domain-specific and general-domain tasks across different models. Results show our method achieves a more favorable balance among inference efficiency, precise length control, and high text quality. Furthermore, unlike previous work, our method efficiently adapts to various length types. In the future, we plan to extend our method to constraints beyond output length, further enhancing LLMs’ controllability over their outputs.
Limitations
This study incorporates length feedback from an external counter to improve LLMs’ ability to regulate output length. Theoretically, this introduces only minor latency due to the efficiency of length computation; however, it interrupts the batched generation process, which can lead to latency issues in large-scale deployment. One promising solution is to selectively decide when to invoke the length counter and to optimize for efficacy xu2025alignment, which we leave for future work.
Appendix A Additional Results from the Preliminary Study
We extend the investigation from token-level to word and sentence granularities. Following the token-level setup, we collect responses of varying lengths by prompting LLMs under the constraint of different target lengths. Specifically, the target lengths at the word level are randomly sampled within the range of 100 to 400 words, whereas sentence‐level target lengths are randomly sampled between 5 and 30 sentences.
Detailed results for the word-level evaluation are shown in Figure 7, and the sentence-level results appear in Figure 8. Consistently across all granularities, longer texts exhibit higher internal estimation errors, indicating a degraded self-assessment capability that impairs precise length control.
Appendix B Analysis of Model Judgment Reliability
To assess the reliability of model-based quality judgments, we conduct a human evaluation study utilizing 50 summarization samples. Human annotators rate each sample along three dimensions namely Coherence, Consistency and Relevance. We then compare these human judgments with the ratings provided by two LLM evaluators, GPT-4.1 and DeepSeek-V3, while measuring the agreement between human and model assessments.
As shown in Table 3, GPT-4.1 achieves a slightly higher consistency ratio for Coherence whereas DeepSeek-V3 outperforms GPT-4.1 in the other two critical dimensions as it attains an 86% consistency ratio for Consistency and a 78% ratio for Relevance. This finding indicates that DeepSeek-V3 possesses greater credibility in capturing human-perceived text quality, leading to its adoption for quality evaluation across all our experiments.
| Coherence | Consistency | Relevance | |
| DeepSeek-V3 | 72% | 86% | 78% |
| GPT-4.1 | 74% | 80% | 68% |
Appendix C Additional Analyses on Training-Free Length Control
We present detailed length distributions on GovReport for Qwen3-4B (Figure 9), Qwen3-8B (Figure 10) and LLaMA-3.1-8B-Instruct (Figure 11). Each figure compares four training-free strategies at the token, word, and sentence granularities across a corresponding range of target lengths. Taking Qwen3-4B as an example, Figure 9(a) to 9(c) contrast Baseline and Feedback, while Figure 9(d) to 9(f) compare ICL and ICL+Feedback. Corresponding evaluations on Biographies appear in Figures 12, 13 and 14. Across various experimental setups, integrating length feedback (i.e., Feedback and ICL+Feedback) consistently yields output length distributions that more closely adhere to target ranges than their counterparts (i.e., Baseline and ICL), especially for larger models. This robust performance underscores the effectiveness and generalizability of the feedback mechanism in achieving precise output length regulation without requiring training.
Appendix D Additional Analyses on Training-Based Length Control
Figure 15 illustrates the length distributions produced by LLaMA-3.1-8B-Instruct with SFT+Feedback and SFT across various target lengths. We observe that LLaMA-3.1-8B-Instruct with SFT+Feedback consistently produces distributions that are more tightly concentrated around the specified target lengths than SFT, indicating superior precision in length control.
Appendix E Hyperparameters for SFT Training
For all models, we perform SFT using the TRL library444https://github.com/huggingface/trl on four A100 GPUs. We set the learning rate to , with a batch size of 64 for training one epoch.
Appendix F Cross-Task Generalization
To verify the model’s cross-task generalization ability after training, we conduct evaluations on TruthfulQA, a dataset unseen during training. As shown in Table 4, our SFT+Feedback model achieves superior length control performance, demonstrating enhanced generalization capability.
| Token | Word | Sentence | |||||||
| MAE | PM | Qual. | MAE | PM | Qual. | MAE | PM | Qual. | |
| SFT | 25.43 | 28.63 | 3.87 | 29.71 | 22.69 | 3.89 | 2.73 | 16.84 | 3.88 |
| SFT+Feedback | 17.83 | 49.32 | 3.86 | 14.52 | 54.96 | 3.82 | 0.52 | 78.10 | 3.76 |
Appendix G Prompt Templates
G.1 Prompt Template for Token Counting
In the preliminary study, we adopt the prompt in Figure 16 to enable LLMs to estimate their output length.
G.2 Prompt Templates for Quality Evaluation
Prompt templates for assessing response quality across various generation tasks are depicted in Figures 17 through 22. In the summarization task, the coherence, consistency and relevance evaluation templates are respectively shown in Figures 18, 19 and 20. For the biography generation task, Figure 21 displays the coherence evaluation template, and Figure 22 presents the template for factuality evaluation. Additionally, Figure 17 shows the quality evaluation template applicable to general-domain text generation. We present all prompt templates employed for text quality evaluation across various generation tasks. Specifically, general-domain texts are measured in terms of correctness and helpfulness. Summaries are evaluated for coherence, consistency, and relevance, while biographies are assessed for coherence and factuality. These prompt templates facilitate a rigorous and uniform approach to measuring key quality dimensions across different generation settings.

### v2

@@
 However, despite significant advancements in following human instructions, Large Language Models (LLMs) still struggle with this task.
-In this work, we demonstrate that LLMs often fail to accurately measure input text length, leading to poor adherence to length constraints.
+In this work, we demonstrate that LLMs often fail to accurately measure their response lengths, leading to poor adherence to length constraints.
 To address this issue, we propose a novel length regulation approach that incorporates dynamic length feedback during generation, enabling adaptive adjustments to meet target lengths.
@@
 1 Introduction
-The use of products built upon Large Language Models (LLMs, achiam2023gpt; team2023gemini; yang2025qwen3) to process various text generation tasks has grown prevalent in daily human study and work.
-Benefiting from pretraining on massive corpora and subsequent post-training, such as Supervised Fine-Tuning (SFT, taori2023alpaca; chiang2023vicuna) and Reinforcement Learning (RL, ouyang2022training; bai2022training), current LLMs demonstrate strong text comprehension and instruction-following capabilities, enabling them to effectively handle diverse user requests.
+The use of products built upon Large Language Models (LLMs, Achiam et al.
+2023; Team et al.
+2025; Yang et al.
+2025) to process various text generation tasks has grown prevalent in daily human study and work.
+Benefiting from pretraining on massive corpora and subsequent post-training, such as Supervised Fine-Tuning (SFT, Taori et al.
+2023; Chiang et al.
+2023) and Reinforcement Learning (RL, Ouyang et al.
+2022; Bai et al.
+2022), current LLMs demonstrate strong text comprehension and instruction-following capabilities, enabling them to effectively handle diverse user requests.
 Despite their significant progress, it remains challenging for LLMs to precisely control their outputs according to specific requirements.
 For instance, when given a query such as “Please write a 500-word composition,” current LLMs often struggle with strictly adhering to the length constraint.
-To address this issue, previous research jie_prompt-based_2024; yuan2024following; li2024ruler; butcher_precise_2025 has trained LLMs to produce outputs that adhere to specified length requirements through SFT or advanced RL algorithms.
+To address this issue, previous research Jie et al.
+(2024); Yuan et al.
+(2024); Li et al.
+(2024); Butcher et al.
+(2025) has trained LLMs to produce outputs that adhere to specified length requirements through SFT or advanced RL algorithms.
 However, these approaches often demand massive training data to cover a wide range of length preferences, and may degrade model performance after intensive post-training.
@@
 Inspired by these findings, we explore the use of length feedback to help LLMs track and dynamically adjust their output length.
-Previous research markergen successfully incorporated word-level length guidance through a three-stage pipeline, inserting length feedback after tokens with an exponential decay strategy.
+Previous research Yuan et al.
+(2025) successfully incorporated word-level length guidance through a three-stage pipeline, inserting length feedback after tokens with an exponential decay strategy.
 However, this method incurs high inference overhead, and its iterative process can lead to error propagation, potentially harming response quality.
@@
 A comparison with the conventional prompt-based approach is shown in Figure 1.
-We first validate our method on summarization huang-etal-2021-efficient and biography generation min2023factscore tasks.
-Without specialized fine-tuning, our method demonstrates superior length control across different constraint types when applied to LLaMA-3.1-8B-Instruct dubey2024llama and Qwen3 yang2025qwen3 family models.
-Evaluations based on DeepSeek-V3 deepseekai2024deepseekv3technicalreport indicate that the text quality remains robust while meeting these constraints.
-To assess our method’s effectiveness on general-domain tasks, we further conduct experiments on ELI5 eli5, a benchmark for long-form question answering.
+We first validate our method on summarization Huang et al.
+(2021) and biography generation Min et al.
+(2023) tasks.
+Without specialized fine-tuning, our method demonstrates superior length control across different constraint types when applied to LLaMA-3.1-8B-Instruct Dubey et al.
+(2024) and Qwen3 Yang et al.
+(2025) family models.
+Evaluations based on DeepSeek-V3 DeepSeek-AI (2024) indicate that the text quality remains robust while meeting these constraints.
+To assess our method’s effectiveness on general-domain tasks, we further conduct experiments on ELI5 Fan et al.
+(2019), a benchmark for long-form question answering.
 While our training-free approach can better meet length requirements, it may compromise quality for tasks requiring answers within a strict length range.
@@
 Data Preparation
-Following prior work jie_prompt-based_2024; butcher_precise_2025, we conduct experiments on the summarization task, where length constraints are commonly specified.
-We use the GovReport dataset huang-etal-2021-efficient and evaluate the following LLMs: LLaMA-3.1-8B-Instruct dubey2024llama, as well as Qwen3-4B and Qwen3-8B yang2025qwen3, chosen for their strong reasoning capabilities on challenging tasks.
+Following prior work Jie et al.
+(2024); Butcher et al.
+(2025), we conduct experiments on the summarization task, where length constraints are commonly specified.
+We use the GovReport dataset Huang et al.
+(2021) and evaluate the following LLMs: LLaMA-3.1-8B-Instruct Dubey et al.
+(2024), as well as Qwen3-4B and Qwen3-8B Yang et al.
+(2025), chosen for their strong reasoning capabilities on challenging tasks.
 To obtain responses of varying lengths, we prompt the LLMs to generate summaries for each instance with a randomly selected target length between 100 and 400 tokens.
@@
 Formally, given an unfinished output , we append a feedback string (e.g., “<used_tokens=>”) before resuming generation, where is a predefined length calculation function.
-Depending on the desired length constraint type (e.g., tokens, words, sentences), can be easily implemented using the model’s tokenizer or external tools like NLTK loper2002nltk.
+Depending on the desired length constraint type (e.g., tokens, words, sentences), can be easily implemented using the model’s tokenizer or external tools like NLTK Bird and Loper (2004).
 Specifically, during generation, we detect sentence boundaries and invoke the length-calculation tool based on the specified constraint type.
@@
 We evaluate our approach across three different models on two tasks suited for length constraints, as well as a general-domain long-form generation task, to assess method robustness.
-GovReport huang-etal-2021-efficient: A long-document summarization dataset of 973 U.S.
+GovReport Huang et al.
+(2021): A long-document summarization dataset of 973 U.S.
 government reports with expert-written abstractive summaries for testing.
-Unlike conventional short-form summarization datasets cohan-etal-2018-discourse; kornilova-eidelman-2019-billsum; sharma-etal-2019-bigpatent, GovReport allows better evaluation across various target lengths.
-Biographies min2023factscore: A biography generation dataset of 183 individuals for testing, originally proposed to evaluate the factuality of LLM-generated text.
+Unlike conventional short-form summarization datasets Cohan et al.
+(2018); Kornilova and Eidelman (2019); Sharma et al.
+(2019), GovReport allows better evaluation across various target lengths.
+Biographies Min et al.
+(2023): A biography generation dataset of 183 individuals for testing, originally proposed to evaluate the factuality of LLM-generated text.
 Unlike summarization, this task relies on LLMs to generate content based on their parametric knowledge.
-ELI5 eli5: A general-domain long-form question-answering dataset.
+ELI5 Fan et al.
+(2019): A general-domain long-form question-answering dataset.
 We use a publicly available dataset111https://huggingface.co/datasets/sentence-transformers/eli5, randomly sampling 10,240 instances for training and 500 for testing.
@@
 We use the same models as in the preliminary study (§2): Qwen3-4B, Qwen3-8B, and LLaMA-3.1-8B-Instruct.
-For all experiments, we conduct inference using the vLLM engine kwon2023efficient, which provides high-throughput serving for LLMs.
+For all experiments, we conduct inference using the vLLM engine Kwon et al.
+(2023), which provides high-throughput serving for LLMs.
 We employ nucleus sampling with temperature set to 0.8 and top-p value of 0.95 for text generation.
 We define target lengths across three linguistic granularities: token-level (100 to 400 tokens, in increments of 50), word-level (100 to 400 words, in increments of 50), and sentence-level (5 to 30 sentences, in increments of 5).
-To quantify generation length across different linguistic units, we use each model’s native tokenizer for token-level counts and the NLTK loper2002nltk for word- and sentence-level measurements.
+To quantify generation length across different linguistic units, we use each model’s native tokenizer for token-level counts and the NLTK Bird and Loper (2004) for word- and sentence-level measurements.
 Evaluation Metrics
-Following prior studies butcher_precise_2025; li2024ruler, we evaluate length‐generation error and text quality at these target lengths using the metrics defined below.
+Following prior studies Butcher et al. (2025); Li et al. (2024), we evaluate length‐generation error and text quality at these target lengths using the metrics defined below.
 Mean Absolute Error (MAE): This metric computes the average absolute difference between the predicted length and the ground-truth user-specified length across all samples:
@@
 (2) where denotes the indicator function. In our experiments, we set for token and word counts, and for sentence-level measurements.
-Quality Score (Qual.): We adopt the LLM-as-a-Judge paradigm zheng2023judgingllmasajudgemtbenchchatbot, using DeepSeek-V3222DeepSeek-V3-0324 across all datasets for more accurate assessments.
-Specifically, for GovReport, we follow liu-etal-2023-g to evaluate summaries on three dimensions: coherence, consistency, and relevance.
+Quality Score (Qual.): We adopt the LLM-as-a-Judge paradigm Zheng et al.
+(2023), using DeepSeek-V3222DeepSeek-V3-0324 across all datasets for more accurate assessments.
+Specifically, for GovReport, we follow Liu et al.
+(2023) to evaluate summaries on three dimensions: coherence, consistency, and relevance.
 For Biographies, we assess coherence and factual accuracy.
-For ELI5, we follow zheng2023judgingllmasajudgemtbenchchatbot by focusing on correctness and helpfulness.
+For ELI5, we follow Zheng et al.
+(2023) by focusing on correctness and helpfulness.
 To manage costs, for each linguistic unit and specified length, we randomly select 100 instances for summary or biography evaluation and 200 instances for general-domain text generation.
@@
 Given the high cost of fine-tuning LLMs, we first compare training-free approaches.
-In addition to conventional prompting with target length constraints (Baseline), we investigate the effectiveness of in-context learning (ICL), which exhibits strong performance across various tasks without task-specific fine-tuning wies2023learnability; dong2024survey.
+In addition to conventional prompting with target length constraints (Baseline), we investigate the effectiveness of in-context learning (ICL), which exhibits strong performance across various tasks without task-specific fine-tuning Wies et al.
+(2023); Dong et al.
+(2024).
 To implement this, we collect instances of varying lengths and construct a one-shot demonstration for each instance.
-Besides, we implement Markgen markergen as a competitive baseline, given that its official code is not publicly available.
+Besides, we implement Markgen Yuan et al.
+(2025) as a competitive baseline, given that its official code is not publicly available.
 Results on GovReport and Biographies are in Table 1. Our key findings are as follows:
@@
 For the feedback-based setting, we first segment responses into sentences, then append length feedback signals after each sentence.
-In addition, we implement RULER li2024ruler, a method that employs a meta-token for word-level length control.
+In addition, we implement RULER Li et al.
+(2024), a method that employs a meta-token for word-level length control.
 We utilize the official code333https://github.com/Geaming2002/Ruler for this implementation.
@@
 Training details are reported in Appendix E.
-Results on ELI5 are presented in Table 2. The following conclusions can be drawn:
 Incorporating length feedback without fine-tuning sometimes degrades performance on the general domain
@@
 Results in Figure 6 demonstrate that SFT+Feedback consistently outperforms the baseline, with a higher density around the target length in each violin plot.
-Furthermore, we assess cross-task generalization on TruthfulQA lin-etal-2022-truthfulqa.
+Furthermore, we assess cross-task generalization on TruthfulQA Lin et al.
+(2022).
 As shown in Appendix F, our model achieves better cross-task generalization.
@@
 Controlling output length is crucial to meet diverse user requirements.
-Early studies addressed this challenge by designing early-stopping decoding strategies rush2015neural or incorporating specialized length-aware module liu2018controlling; takase2019positional; yu2021lenatten to guide generation.
-With the emergence of LLMs achiam2023gpt, their advanced instruction-following capability enables length control through direct prompting markergen.
-However, existing methods still rely on extensive post-training, such as SFT yuan2024following; li2024ruler; butcher_precise_2025 or RL singhal2023long; jie_prompt-based_2024, to achieve precise length control.
+Early studies addressed this challenge by designing early-stopping decoding strategies Rush et al.
+(2015) or incorporating specialized length-aware module Liu et al.
+(2018); Takase and Okazaki (2019); Yu et al.
+(2021) to guide generation.
+With the emergence of LLMs Achiam et al.
+(2023), their advanced instruction-following capability enables length control through direct prompting Yuan et al.
+(2025).
+However, existing methods still rely on extensive post-training, such as SFT Yuan et al.
+(2024); Li et al.
+(2024); Butcher et al.
+(2025) or RL Singhal et al.
+(2024); Jie et al.
+(2024), to achieve precise length control.
 In contrast, our method accomplishes this goal by incorporating length feedback, which is either training-free or training-efficient.
@@
 5.2 Feedback-Guided Text Generation
-Since LLMs are essentially probabilistic models, they often struggle with tasks involving complex mathematical reasoning zhou2022least or up-to-date factual knowledge luo2023systematic.
-A straightforward way to overcome these limitations is to equip LLMs with the ability to use external tools parisi2022talm; schick2023toolformer, such as search engines, calculators, or compilers.
+Since LLMs are essentially probabilistic models, they often struggle with tasks involving complex mathematical reasoning Zhou et al.
+(2023) or up-to-date factual knowledge Luo et al.
+(2023).
+A straightforward way to overcome these limitations is to equip LLMs with the ability to use external tools Parisi et al.
+(2022); Schick et al.
+(2023), such as search engines, calculators, or compilers.
 Using external feedback from these tools, LLMs’ capabilities can be substantially extended, allowing them to focus better on tasks they inherently excel at.
-Markergen markergen was the first to integrate length feedback directly into the generation process.
+Markergen Yuan et al.
+(2025) was the first to integrate length feedback directly into the generation process.
 However, its multi-stage framework introduces considerable complexity and inefficiency.
@@
 Theoretically, this introduces only minor latency due to the efficiency of length computation; however, it interrupts the batched generation process, which can lead to latency issues in large-scale deployment.
-One promising solution is to selectively decide when to invoke the length counter and to optimize for efficacy xu2025alignment, which we leave for future work.
+One promising solution is to selectively decide when to invoke the length counter and to optimize for efficacy Xu et al.
+(2025), which we leave for future work.
 Appendix A Additional Results from the Preliminary Study
@@
 To assess the reliability of model-based quality judgments, we conduct a human evaluation study utilizing 50 summarization samples.
-Human annotators rate each sample along three dimensions namely Coherence, Consistency and Relevance.
-We then compare these human judgments with the ratings provided by two LLM evaluators, GPT-4.1 and DeepSeek-V3, while measuring the agreement between human and model assessments.
+We recruited five graduate students who verified their fluency in English and familiarity with fundamental text evaluation concepts.
+Each annotator was compensated at a flat rate of $15 per hour.
+To ensure uniform understanding of the comparative evaluation criteria, all annotators received detailed written instructions and participated in a brief training session.
+The full text of the unified instructions is provided in Figure 9.
+For each sample, human annotators are presented with two summaries: one generated by the feedback-based method and one generated by the baseline method.
+The annotators are blinded to the source of each summary.
+Their task is to compare the two summaries along three dimensions: Coherence, Consistency, and Relevance.
+We compare these human comparative judgments with the ratings provided by two LLM evaluators, GPT-4.1 and DeepSeek-V3, respectively.
+Both models assign individual scores to each of the two summaries separately, a practice that generates comparative signals which enable alignment with the comparative judgments derived from human evaluators.
 As shown in Table 3, GPT-4.1 achieves a slightly higher consistency ratio for Coherence whereas DeepSeek-V3 outperforms GPT-4.1 in the other two critical dimensions as it attains an 86% consistency ratio for Consistency and a 78% ratio for Relevance.
@@
 Each figure compares four training-free strategies at the token, word, and sentence granularities across a corresponding range of target lengths.
-Taking Qwen3-4B as an example, Figure 9(a) to 9(c) contrast Baseline and Feedback, while Figure 9(d) to 9(f) compare ICL and ICL+Feedback.
-Corresponding evaluations on Biographies appear in Figures 12, 13 and 14.
+Taking Qwen3-4B as an example, Figure 10(a) to 10(c) contrast Baseline and Feedback, while Figure 10(d) to 10(f) compare ICL and ICL+Feedback.
+Corresponding evaluations on Biographies appear in Figures 13, 14 and 15.
 Across various experimental setups, integrating length feedback (i.e., Feedback and ICL+Feedback) consistently yields output length distributions that more closely adhere to target ranges than their counterparts (i.e., Baseline and ICL), especially for larger models.
@@
 G.2 Prompt Templates for Quality Evaluation
-Prompt templates for assessing response quality across various generation tasks are depicted in Figures 17 through 22.
-In the summarization task, the coherence, consistency and relevance evaluation templates are respectively shown in Figures 18, 19 and 20.
+Prompt templates for assessing response quality across various generation tasks are depicted in Figures 18 through 23.
+In the summarization task, the coherence, consistency and relevance evaluation templates are respectively shown in Figures 19, 20 and 21.
 For the biography generation task, Figure 22 displays the coherence evaluation template, and Figure 23 presents the template for factuality evaluation.